Many species have a light and dark form. The dark form is usually the dominant allele. For peppered moths, the light form provides camouflage on trees covered with lichen, protecting them from bird predation. The dark form was only present at about a rate of 1% in the population. With the onset of the Industrial Revolution in the 1850’s, the frequency of the dark form increased to about 90% as the dark moths could camouflage better on soot covered trees. Dead light and dark moths were pinned to light or dark backgrounds and the number taken by birds recorded. Genotype DD Dd dd Phenotype dark dark light Fitness on light background 0.65 0.65 1.0 Fitness on dark background 1.0 1.0 0.56 What is the equilibrium frequency of D in the presence of light background? In the presence of the dark background? What is the mean relative fitness for light and dark?
Answer #1

Equilibrium frequency of D(dark form allele) in the presence of light background can be calculated as

q =s1/(s1+s2) where s1 and s2 are the selection coefficients of the two homozygotes DD and dd.

Here s1 = (1-0.65) = 0.35 and s2 = (1-1) =0

Therefore, q = 0.35/(0.35+0) = 1

Equilibrium frequency of D(dark form allele) in the presence of light background is 1.

Equilibrium frequency of D(dark form allele) in the presence of dark background can be calculated as

q =s1/(s1+s2) where s1 and s2 are the selection coefficients of the two homozygotes DD and dd.

Here s1 = (1-1) = 0 and s2 = (1-0.56) =0.44

Therefore, q = 0/(0+0.44) = 0

Equilibrium frequency of D(dark form allele) in the presence of dark background is 0.

Mean relative fitness for light background can be calculated as

W = 1-sq2 = 1-(0.35*1*1) = 1-0.35 = 0.65

Mean relative fitness for dark background can be calculated as

W = 1-sq2 = 1-(0.44*0*0) = 1-0 = 1